a stone is thrown in such a manner that it would just hit a bird at the top of a tree and afterwards reach a maximum height double that of a tree .if at any moment of throwing the stone the bird flies away horizontally with constant velocity and the stone hits the bird after some time .the ratio of horizontal velocity of stone to that of bird is 1/n+1/root x .find 2n

Grade:11

1 Answers

Divya shukla
6 years ago
t1 = √(2h/g) * [√2 - 1] t2 = √(2 h/g) * [√2 + 1 ]at t = t1, the stone travels a distance = ux * t1at t = t2, the stone travels a distance = ux * t2 = ux * t1 + v * t2so ux / v = t2/(t2 - t1) = (√2 + 1)/2 = 1/√2 + 1/2So n = 2 and 2n = 4